OpenStudy (anonymous):

hello, How can i convert from sine/cosine/tangnet to radians? ex: arcsin(1/2) Can someone go with me step by step? thanks.

OpenStudy (anonymous):

We know that one of the angles \[\theta\] such that \[\sin\theta=\frac12\] is : \[\theta =30^\circ\] So we should first convert it to radians like this : \[30^\circ=\frac{30\pi}{180}=\frac\pi6\text{rd}\] So : \[\arcsin \frac12=\frac\pi6\]